Wednesday 16 3
This presentation is the property of its rightful owner.
Sponsored Links
1 / 8

Wednesday 16/3 PowerPoint PPT Presentation


  • 47 Views
  • Uploaded on
  • Presentation posted in: General

Wednesday 16/3. 08:30 – Dumped physics fill with 3 x 3 bunches AM: Investigations on tune feedback Ramped 8 nominal bunches again, which stayed in this time, although perturbations still there. Collide 8 bunches. Dump. PM: No beam

Download Presentation

Wednesday 16/3

An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -

Presentation Transcript


Wednesday 16 3

Wednesday 16/3

  • 08:30 – Dumped physics fill with 3 x 3 bunches

  • AM: Investigations on tune feedbackRamped 8 nominal bunches again, which stayed in this time, although perturbations still there. Collide 8 bunches. Dump.

  • PM: No beam

    • Problems with RF re-phasing: “there is a large static error on the synchro loop”

    • BST MASTER reboot

    • PS no beam

  • Evening: Continue investigations.Spectrum looks very different for 32 bunches compared to 8 bunches

    • 20:40 Ramp 32 nominal bunches

    • Lost because of RQTD trips – but different signature

    • Problem with timing distribution

  • Night: Requalification of systems. Collimator problem.

LHC 8:30 meeting


The 8 bunch ramp that stayed in lucky

The 8 bunch ramp that stayed in - lucky

Tune-trim B2 H

OK when tunes moved for collisions

LHC 8:30 meeting


Tune feedback spectrum issue

Tune feedback / spectrum issue

Reduction of damper gain

  • Problem:With 8 nominal bunches beams were lost at 3.5 TeV twice on Tuesday. “100 Hz” peaks in tune spectrum for B2 H makes tune feedback jump around, resulting in RQT trip.

  • At injection: Changing damper gains at injection -> makes tune measurement more stable

  • At 3.5 TeV switch off ADT -> perturbing lines still too large relative to tune lines.

  • LHC 8:30 meeting


    32 bunches at injection spectra look different

    32 bunches at injection – spectra look different

    32 bunches

    8 bunches

    ADT off

    LHC 8:30 meeting


    Ramp 32 bunches

    Ramp 32 bunches

    Jump in Time

    • Lost in ramp RQTD.A12B2 and RQTD.A67B2 tripped off

    • But tune signals were proper this time

    • Previous trips were RQTF and at flat top

    • RQTD U-res not noisy, but is sudden spike

    • PM, XPOC, logging thoroughly messed up

      • Switching the LHC Timing Gateway from A->B

    LHC 8:30 meeting


    Qps of rqtd reacted correctly

    QPS of RQTD reacted correctly

    LHC 8:30 meeting


    Recover

    Recover

    • As a result of this timing card switch, preformed a set of checks to confirm that the timing was OK and that the dump and PM was behaving as expected. Tests done:

      • Dry dump to check PM

      • Dump with pilot at injection to test PM etc

      • Ramp with pilots and then taken through the collisions BP

    • Tried to fill physics with 3 bunches – waiting for experts to report on “timing issue”

      • Collimation problem, TCTH.4L5.B1 position interlock at start of squeeze, expert working.

      • BPM calibration problem

    LHC 8:30 meeting


    The plan

    The Plan

    • Confirmation from experts everything ok

      • Timing

      • Collimation

    • Ramp 32 bunches, aiming for stable beams

    • To do:

      • Collimation energy interlocks

    LHC 8:30 meeting


  • Login